The ratio of boys to girls in a school is 3 : 4 .
There are 120 more girls than boys.
How many students are in the school?

Answer by stanbon(75887)

The ratio of boys to girls in a school is 3 : 4 .
There are 120 more girls than boys.
How many students are in the school?
-------
3:4 is the same as 3x:4x where 3x is the # of boys: 4x is the # of girls
--------
Equation:
4x = 3x+120
x = 120
----
Ans: 3x+4x = 7x = 7*120 = 840
-----
Cheers,
Stan H.
------